The Function 'f' is defined by f(x)=ax2+bx+c.Given that f(0)=6,f(-1)=15 and f(1)=1,find the values of a,b,c
Answer by dkala99(12)

f(x)=ax^2+bx+c
f(0) = 6
a(0)^2+b(0)+c = 6
0+0+c = 6
c = 6
f(-1) = 15
a(-1)^2+b(-1)+c = 15
a-b+c = 15
a-b+6 = 15
a-b = 15-6
a-b = 9 -----> equ 1
f(1) = 1
a(1)^2+b(1)+c = 1
a+b+c = 1
a+b+6 = 1
a+b = 1-6
a+b = -5 -----> equ 2
solve equ 1 and 2
a-b = 9 (+)
a+b = -5
---------
2a = 4
a = 2
from equ 2
a+b = -5
2+b = -5
b = -5-2
b = -7
a = 2 ; b = -7 ; c = 6
-----------------------------------
check in f(1)=1
f(1) = 2(1)^2+(-7)(1)+6
= 2-7+6
f(1)= 1